# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 2.5 Operational Design Domains and Micro ODDs
Assurance goals for a system can be lessened by limiting the range or complexity of circumstances (i.e., environments) in which it is required to operate. For example, self-driving is easier on freeways or in traffic jams than on city streets. These different circumstances are referred to as Operational Design Domains (ODDs) and a system may be assured only for specific ODDs and be required to disengage when outside those permitted (alternatively, the system may have different modes of operation in different ODDs). Clearly, the perception system must be augmented to determine when it is in permitted or specific ODDs.
@@
A further enhancement might be to use multiple backup guards, each specialized to a particular circumstance or ODD. The idea is that forward-looking radar coupled to emergency braking may provide an assured backup arrangement suitable for driving in traffic, but for highway driving it would be better to look further ahead using cameras (having non-AI perception) with speed control as the intervention. These ODDs would be tailored to assured detection and intervention strategies and might not correspond to traditional ODDs: following [99] we call them micro-ODDs (also written as $\mu$ODDs). The idea is that for any particular micro-ODD, using the “right” backup guard will deliver superior safety, with fewer false alarms and less disruptive interventions.
This suggests an architecture such as portrayed in (8) below where a portfolio of assured backup guards is coordinated by a detector that recognizes the current micro-ODD and selects the appropriate assured backup guard. Of course, the perception system of the detector must be assured, but its task seems to be rather simple and it is conceivable that it can be performed without AI, or by AI and ML of credible trustworthiness (see [115] for relevant work).
@@
    Architecture (7) with a portfolio of assured backup guards coordinated by an assured detector that recognizes their matching micro-ODDs.
@@ 3 Assurance of AI Systems for Specific Functions
In this section we consider novel non-CPS systems and applications that are made possible or are performed in new ways by the capabilities of AI and ML. These include systems that play games of skill or strategy, those that design things or perform scientific predictions such as protein folding or weather forecasting, the generation or management of responses based on these, decision support systems that analyze medical images or loan, job, and college applications or prisoner sentencing and parole and so on, and also systems that use Large Language Models (LLMs) such as ChatGPT and other general-purpose AI and ML capabilities to perform specific functions, which can include generation of images and video as well as text. We do not include general-purpose capabilities themselves (those are in the next section).
@@
The reason we focus here on systems that perform specific functions is that assurance needs to identify hazards; the hazards of systems that perform specific functions can be conjectured from the functions concerned and the environment in which they are deployed, whereas the hazards of general-purpose systems such as LLMs have to consider all the functions they might be called to perform and all the environments they might be deployed in.
@@
Hazards of specific applications include: making incorrect or poor decisions, generation or approval of offensive or untrue material, exhibiting bias or stereotyping in any of these activities, causing distress, enabling crime (e.g., extortion using voice clones), vulnerability to manipulation, and so on. Many of these are quite different to the hazards of traditional systems, so standard methods of hazard analysis may be difficult to apply. However, hazards specific to AI and ML can often be anticipated by considering poor or malicious human performance and interaction in a similar context. For example, Microsoft’s “Tay” was a Twitter bot that the company described as an experiment in “conversational understanding.” The more you chat with Tay, said Microsoft, the smarter it gets, learning to engage people through “casual and playful conversation.” Within less than a day of its release, it had been trained by a cadre of bad actors to behave as a racist mouthpiece and had to be shut down [178]. This is a hazard that should have been anticipated.
@@
Similar to the detection of hazards, tolerable failure rates may also be estimated by comparison with human performance, although society may be less tolerant of failures by an AI system than those by humans (recall the discussion on fatality rates for self-driving cars on page 1.2).
@@
We continue to refer to assured hazard elimination generically as “dependability” even though the specific hazards may not concern risk to life or conventional assets. Assurance can then be founded on similar principles as it is for traditional systems, and required confidence will be graduated according to severity of the hazards [44]. Dependability requirements validation, the first step in traditional assurance, can proceed for an AI application rather as it does for a traditional system, or for a CPS system augmented with AI (recall Sections 1 and 2). That is, hazards should be identified, requirements should be developed to eliminate or mitigate them, and analysis should demonstrate that they do so. Some hazards and their controls may conflict: for example, control of offensive material may conflict with free speech, and suitable policies and compromises must be developed. But these conflicts exist in human systems, it is just that normally we do not have to document explicit choices as we do when formulating dependability requirements (consider recent difficulties in university responses to student protests).
@@
While it seems plausible that dependability requirements validation can be performed for systems powered by AI and ML much as it is for conventional systems, differences and difficulties arise at the next stage. These are the verification tasks where, as we saw in the previous section, specifications are likely to be absent and there are no strong reasons to justify intent or correctness verification other than statistical observation. However, unlike the safety-critical applications considered in the previous section, those contemplated here may require only modest levels of confidence with respect to dependability (because they do not pose immediate high-rate hazards to life or critical assets) and it is possible that statistically valid testing could contribute adequate assurance, particularly when coupled (via CBI) with methods for prior confidence such as careful selection of training data, and architectural mechanisms for mitigation such as diverse redundancy and runtime checking.
@@
For example, if racial bias is recognized as a hazard, then it might be mitigated by removing race from the data presented to the ML in training and operation. A weakness in this approach is that the ML may discover a proxy for race (e.g., zip code) among the data that it does see, so a better alternative may be to mask this characteristic in training by assigning race randomly.
@@
More likely, however, assurance must depend on some form of runtime verification, implemented either by additional training and careful prompting for “fine tuning” the LLM,191919 LLMs are adapted to specific tasks by instructions, referred to as “prompts,” given in natural language [127]. or by explicit guards. As with the autonomous systems of the previous section, runtime verification poses difficulties due to lack of non-AI ways to perceive the context or world model within which to make the guarding decisions. For example, if an LLM is suspected of racial bias despite race being absent from its training and operational data, a guard also lacking true data on race might need to have its own ML component to “perceive” this attribute.
@@
It is worth taking a short detour here to examine the general unreliability of LLMs. In their “chatbot” manifestations (and hence, by extension, in specific applications), it is well-known that LLMs can generate fluent but false or meaningless utterances. These are often referred to as “hallucinations” [139] although some critics prefer other terms, such as “fabrications” [137]. In our opinion, these and similar terms are inappropriate as they suggest the LLM has some (mis)perception of the world and awareness of truth and falsehood. In reality, LLMs are trained simply to predict the next or missing “token” in a string of text based on statistical observations of a vast corpus; in one memorable phrase, they are “stochastic parrots” [14]. We will use the neutral term “failure” to refer to all kinds of false, offensive, or unhelpful responses.
@@
Effective cooperation, and communication in particular, are based on the parties having a shared context or world model [55] (as theologians say: “text without context is pretext”). Popper proposed an ontology of Three Worlds [132, 133] that is somewhat controversial, but which finds application in Computer Science [160], and is useful for our purpose: World 1 comprises objects and properties of the physical universe such as those addressed by scientific theories (e.g., planets, mass, motion); World 2 is mental states and processes, or what I (and you) are thinking about; and World 3 is the “products of thought” such as tables and chairs [^20] The physical attributes of tables and chairs are in World 1, but their functions (i.e., chairs are for sitting on) are “products of thought” and belong to World 3. and the UK Highway (driving) Code. Human communication relies on shared models for a selection of these that are relevant to the current conversation. In particular, I need approximate models of some of your models: for example, if I am your driving instructor, I need a World 2 model of your World 3 model of the Highway Code. An LLM has none of this: its utterances are model-free and align with the three worlds of any specific context purely by chance and statistical association. Hence, the model-free utterances of LLMs are simply unconnected with the conversational context [72] and frequent failures are to be expected.
@@
On the other hand, LLMs are popular because their performance goes beyond that suggested by “predicting the next or missing ‘token’ in a string of text”: there seems to be emergent behavior that delivers more value than this. Similarly, although there are no explicit models providing context for this behavior, it is possible that implicit models emerge from statistical associations in LLM training data, [^21] For example, recent work finds “features” within LLMs that correspond to concepts and bias [166], and also “correspondence” between representation of language in LLMs and the brain [169]. and this might partially account for their surprising performance. It would also explain their flaws and, further, indicate that these flaws are inevitable [179], unpredictable, and unfixable—unless assurable world models can be incorporated within LLMs, or within explicit guards.
There are many proposals for constructing “guardrails” within or around LLMs. We discuss some of these in the next section on assurance for general-purpose AI but here we will consider some that can be customized to specific applications.
@@
One approach is to develop “workflows” (i.e., sequences of prompts) where the LLM is iteratively asked to critique and improve its previous outputs. A recent paper reports precision around 90% for a workflow that extracts data (as Material, Value, Unit triplets) from materials science papers [131] and similar approaches have been successful in other applications [108]. Another direction exploits the large “context window” (i.e., input) allowed by some recent LLMs to provide a prompt with hundreds of training examples prior to the real query [2] (the paper also proposes ways to automatically generate suitable examples). This is called “in-context learning” [34]; previously, such “fine tuning” required access to the LLM’s training environment and made adjustment to the weights in its neural net. Related to this are applications that provide a substantial input and then ask the LLM to do something with it (e.g., summarize it, or identify its topic).
@@
All these approaches exploit the natural language capability of the LLM, but constrain it to operate on or within the input provided, so there is little opportunity to generate or insert content extracted from its training corpus. Due to its model-free nature, the LLM may still misinterpret the input and do a wrong thing, but this should also be minimized as a large input can explicitly (via extensive prompts) or implicitly (via a block of text) convey the intended context or model.
@@
Another approach operates by attempting to access the model or “reasoning” purportedly employed. This is Explainable AI, which aims to deliver reasons for accepting the output from an LLM. Counterfactual explanations, where a system may deliver a response such as “I am declining your loan application but would have approved it if your income was \$5,000 greater, or your deposit was \$2,000 more” [173, 47], may be particularly useful for specific applications as they should be expressed in terms of the policy specified in the prompt and not arbitrary “facts” from the training corpus. A runtime checker could reject decisions whose counterfactual explanations violate the policy. Do note, however, that counterfactual explanations can be manipulated [159], so this approach needs to be employed with caution. However, in some applications it does not matter how a decision and its explanation are produced: they can be considered valid as long as they pass an independent check (that does incorporate models of the appropriate policy).
@@
A related approach can be applied to AI systems that generate some sort of “design,” such as plans, schedules, software code, or physical artifacts. Even without an explanation, these can often be guarded by analysis tools or simulators for the domain concerned. Again, it does not matter how the proposed design is generated: if it satisfies traditionally engineered and assured verification tools, then it can be considered good (e.g., see [84] for an elementary application to planning).
@@
Automated formalization (or “autoformalization”) is a popular application of this kind that provides an example. The idea is that we have some informal natural language description of the requirements for a computer system, or for a simulation model or for some program code, and the goal is to translate it into a “formal” representation that can be ingested and analyzed, simulated, or executed by some automated verification tool that acts as a guard. A typical approach invites the LLM to generate a formal representation, then uses the verification tool to analyze the result; if this is unsatisfactory, the counterexample, diagnosis, or error report produced by the verification tool is given to the LLM, which is then asked to provide a better formalization. The process iterates until the formal representation passes scrutiny by the verification tool. (This is a form of “CounterExample-Guided Abstraction Refinement” or CEGAR [43].)
@@
Difficulties may arise with this general approach if the AI system proposes designs that are so original they are outside the capabilities of the analysis or simulation performed by the guarding tool (e.g., a laterally asymmetric airplane may defeat an aerodynamic simulator). A related difficulty arises in domains where there is no reliable means of analysis. An example is generation of strategy for games or real-world scenarios. Applications that generate strategies typically use techniques based on reinforcement learning rather than LLMs. One popular approach, related to mechanisms found in all vertebrate brains [18, Section 6] uses an actor and a critic. The actor generates behaviors and the critic predicts the likelihood that each will lead to a successful outcome. The actor is rewarded by favorable predictions and the critic by their accuracy. Starting with random behaviors and predictions plus some tactic for exploration, the pair will generally converge on optimal behavior, given sufficiently many training examples [88]. This provides assurance for examples in the training set but, as always with ML, tells us little about performance on new examples. Here, the only automated way to check plausibility prior to commitment may be to use another (diverse) AI system.

### v3

@@ 2.6 Operational Design Domains and Micro ODDs
Assurance goals for a system can be lessened by limiting the range or complexity of circumstances (i.e., environments) in which it is required to operate. For example, self-driving is easier on freeways or in traffic jams than on city streets. These different circumstances are referred to as Operational Design Domains (ODDs) and a system may be assured only for specific ODDs and be required to disengage when outside those permitted (alternatively, the system may have different modes of operation in different ODDs). Clearly, the perception system must be augmented to determine when it is in permitted or specific ODDs and this determination must be assured to operate with sufficient accuracy.
@@
A further enhancement might be to use multiple backup guards, each specialized to a particular circumstance or ODD. The idea is that forward-looking radar coupled to emergency braking may provide an assured backup arrangement suitable for driving in traffic, but for highway driving it would be better to look further ahead using cameras (with non-AI perception) with speed control as the intervention. These ODDs would be tailored to assured detection and intervention strategies and might not correspond to traditional ODDs: following [140] we call them micro-ODDs (also written as $\mu$ODDs). The idea is that for any particular micro-ODD, using the appropriate backup guard will deliver superior safety, with fewer false alarms and less disruptive interventions.
This suggests an architecture such as portrayed in (8) below, where a portfolio of assured backup guards is coordinated by a detector that recognizes the current micro-ODD and selects the appropriate assured backup guard. Of course, the perception system of the detector must be assured, but its task seems to be rather simple and it is conceivable that it can be performed without AI, or by AI and ML of credible trustworthiness (see [167] for relevant work).
@@
    Architecture 7 with a portfolio of assured backup guards coordinated by an assured detector that recognizes their matching micro-ODDs (and also provides this information to the local model).
Having discussed assurance for systems extended with AI and ML, the attendant problem of assuring perception, and the possibility of assurance through runtime monitoring, diversity, and defense in depth, we next apply these ideas to other AI systems that perform specific functions.
@@ 3 Assurance of AI Systems for Specific Functions
In this section we consider novel non-CPS systems and applications that are made possible or are performed in new ways by the capabilities of AI and ML. These include systems that play games of skill or strategy; those that design things or perform scientific predictions such as protein folding or weather forecasting and the generation or management of responses based on these; decision support systems that analyze medical images or loan, job, and college applications or prisoner sentencing and parole and so on; autonomous “AI Agents” that perform tasks such as travel arrangements and other secretarial functions, and also systems that use Large Language Models (LLMs) such as ChatGPT and other general-purpose AI and ML capabilities to perform these and other specific functions, which can include generation of software, images, speech and video, as well as text. We do not include general-purpose capabilities themselves (those are in the next section).
3.1 Feasibility of Hazard Analysis
@@
The reason we focus here on systems that perform specific functions is that assurance needs to identify hazards; the hazards of systems that perform specific functions can be anticipated or predicted from the functions concerned and the environment in which they are deployed, whereas the hazards of general-purpose systems such as LLMs have to consider all the functions they might be called to perform and all the environments they might be deployed in. But note that while systems for specific functions introduce specific hazards, they may also protect the underlying AI mechanism from other hazards. For example, a system for medical advice has no need, and should be unable, to ask its LLM for instructions on making chemical weapons or hacking software, and this can mitigate one class of risks.
@@
Hazards of specific applications include: making incorrect or poor decisions, generation or approval of offensive or untrue material, exhibiting bias or stereotyping in any of these activities, causing distress, enabling crime (e.g., extortion using voice clones), vulnerability to manipulation, and so on [241]. Many of these are quite different to the hazards of traditional systems, so standard methods of hazard analysis can be difficult to apply and research may be needed to develop suitable new methods. However, hazards specific to AI and ML can often be anticipated by considering poor or malicious human performance and interaction in a similar context. For example, Microsoft’s “Tay” was a Twitter bot that the company described as an experiment in “conversational understanding.” The more you chat with Tay, said Microsoft, the smarter it gets, learning to engage people through “casual and playful conversation.” Within less than a day of its release, it had been trained by a cadre of bad actors to behave as a racist mouthpiece and had to be shut down [245]. This is a hazard that should have been anticipated.
@@
Similar to the detection of hazards, tolerable failure rates may also be estimated by comparison with human performance, although society may be less tolerant of failures by an AI system than those by humans (recall the discussion on fatality rates for self-driving cars on page 2).
@@
We continue to refer to assured hazard elimination generically as “dependability” even though the specific hazards may not concern risk to life or conventional assets. Assurance can then be founded on similar principles as it is for traditional systems, and required confidence will be graduated according to severity of the hazards [62]. Dependability requirements validation, the first step in traditional assurance, can proceed for an AI application rather as it does for a traditional system, or for a CPS system augmented with AI (recall Sections 1.1 and 2). That is, hazards should be identified, requirements should be developed to eliminate or mitigate them, and analysis should demonstrate that they do so. Some hazards and their controls may conflict: for example, control of offensive material may conflict with free speech, and suitable policies and compromises must be developed. But these conflicts exist in human systems, it is just that normally we do not have to document explicit choices as we do when formulating dependability requirements (consider recent difficulties in university policies regarding student protests about the war in Gaza).
Although safety is a system property, implementations that employ AI and ML sometimes focus on these and their trustworthiness rather than the overall system. This compromises hazard analysis and assurance as we now illustrate.
  • Hazards due to causes other than unreliability may not be considered: for example, harms due to malicious use of the AI and the impacts of misalignment.
  • Evaluation of hazards may fail to consider the many pathways to harm, and criticality assessment may be unrealistic and not matched to the assurance that is proposed or feasible. For example, phrases such as “existential risk” suggest the need for safety assurance more rigorous than aircraft software but are used in conjunction with weak means of assurance such as testing and “red teams” and without analyzing the pathways to harm.
  • Often there are no explicit functional requirements, so we do not have a basis for assurance.
Despite these common lapses, it does seem feasible to perform realistic hazard analysis for specific systems that employ AI and ML, although this may require development of new methods or revisions to existing methods as AI and ML may introduce new hazards, and new failure modes might lead to combinations, or increased likelihood, of hazards than were previously considered. It should then be feasible to develop and validate requirements to eliminate or mitigate those hazards, much as is done for conventional systems. However, differences and difficulties arise at the next stage, which includes the verification and assurance tasks.
3.2 Verification and Assurance
@@
As we saw in the previous section, specifications are likely to be absent for AI and ML elements and there are no strong reasons to justify intent or correctness verification other than statistical observation. However, unlike the safety-critical applications considered in the previous section, some of those contemplated here may require only modest levels of confidence in their dependability (because they do not pose immediate high-rate hazards to life or critical assets) and it is possible that a case based on statistically valid testing could deliver adequate assurance providing the underlying assumptions of the statistical approach, including those on the environment and behaviors of the software, are justified. Note that typical methods for benchmarking and testing AI systems do not amount to statistically valid testing and cannot deliver assurance suitable for this purpose.
@@
Unfortunately, it can be difficult to provide assurance that establishes the prior confidence needed for CBI. The root of the difficulty is that we have limited control over the local and domain models built by the ML system and little ability to inspect or review them. Nonetheless, we might hope that careful selection of training data provides some assurable control over the models generated. For example, if racial bias is recognized as a hazard then it might be mitigated by removing race from the data presented to the ML in training and operation. The hope is that by eliminating race from the training data, it will be absent or neutral in the models that ML constructs. However, the ML may discover a proxy for race (e.g., zip code) among the data that it does see, so a better alternative may be to mask this characteristic in training by assigning race randomly.
@@
More likely, however, assurance must depend on some form of runtime verification, implemented either by additional training for “fine tuning” the LLM, or by careful prompting, [^20] LLMs are adapted to specific tasks by instructions, referred to as “prompts,” given in natural language [182]. or by explicit guards. As with the autonomous systems of the previous section, runtime verification poses difficulties due to lack of non-AI ways to perceive the context or local model within which to make the guarding decisions. For example, if an LLM is suspected of racial bias despite race being absent from its training and operational data, a guard also lacking operational data on race might need to have its own ML component to “perceive” this attribute.
@@
It is worth taking a short detour here to examine the general unreliability of LLMs. In their “chatbot” manifestations (and hence, by extension, in specific applications), it is well-known that LLMs can generate or interpolate fluent but false or meaningless utterances. For example, OpenAI’s “Whisper” speech-to-text system can add gratuitous phrases to its transcriptions; in medical record keeping, this can produce high-consequence faults such as incorrect lists of prescribed drugs [135]. The underlying errors are often referred to as “hallucinations” [196] although some critics prefer other terms, such as “confabulations” or “fabrications” [194]. In our opinion, these and similar terms are inappropriate as they suggest the LLM has some understanding of the world and awareness of truth and falsehood. In reality, LLMs are trained simply to predict the next or missing “token” in a string of text based on statistical observations of a vast corpus; in one memorable phrase, they are “stochastic parrots” [19], although this may understate the performance of recent systems. We will use the neutral term “failure” to refer to all kinds of false, offensive, or unhelpful responses.
Like any other system that interacts with the world, an LLM must build a model of its local environment and use it to generate some useful response. Humans likewise build models of the world in order to base their behavior on predicted outcomes [64, 121]. For substantive dialog and cooperation with humans, an LLM must have a model that matches some aspects of the human “mental model,” so that both parties have a shared context [76].
A concrete example is use of an LLM to enable human-robot coordination [93]. The human specifies what they want done, the robot constructs a plan and describes it to the human, who approves it; the robot then executes the plan, describing its progress at each step. It is obvious that this is safe (not to mention feasible) only if the human and robot have very similar models of the environment.
@@
Unfortunately, the structure of human mental models is unclear. Popper proposed they are based on an ontology of Three Worlds [188, 189] that is somewhat controversial, but which finds application in Computer Science [223], and is useful for our purpose: World 1 comprises objects and properties of the physical universe such as those addressed by scientific theories (e.g., mass, motion, the planets); World 2 is mental states and processes, or what I (and you) are thinking about; and World 3 is the “products of thought” such as tables and chairs [^21] The physical attributes of tables and chairs are in World 1, but their functions (i.e., chairs are for sitting on) are “products of thought” and belong to World 3. and the UK Highway (driving) Code. Human communication relies on shared models for a selection from each of these that are relevant to the current dialog. In particular, I need approximate models of some of your models: for example, if I am your driving instructor, I need a World 2 model of your World 3 model of the Highway Code. A significant point is that only World 1 is detected by conventional sensors, yet most of our interactions are with Worlds 2 and 3. Furthermore, World 3 requires substantial domain models: to interact with your car, you need at least a rough idea (you can look up the details) of what it is for, what is does, and how it operates.
An LLM has none of this in explicit form: its utterances are model-free and align with the three worlds of any specific context purely by statistical association. Hence, the utterances of LLMs are simply unconnected with the way the world works or the current conversational context [100] and frequent failures are to be expected. [^22] It may seem perverse to declare that LLMs are model-free when “model” appears in their very name. However, the “model” in LLM refers to an algorithmic model of natural language, whereas we are speaking of models as representations of the system’s “world” or context.
@@
On the other hand, LLMs are popular because their performance goes beyond that suggested by “predicting the next or missing ‘token’ in a string of text”: there seems to be emergent behavior that delivers more value than this. Similarly, although there are no explicit models providing context for this behavior, it is possible that implicit models emerge from statistical associations in LLM training data: the mental models of humans underlie their speech and behavior and it is plausible that LLMs infer at least the linguistic aspects of human mental models and this might partially account for their surprising performance (e.g., recent work finds “correspondence” between representation of language in LLMs and the brain [233] and also “features” within LLMs that correspond to concepts and bias [230]; however, more recent work finds significent difference between the concepts constructed by humans and by LLMs [219]). It would also explain their flaws and, further, indicate that these flaws are inevitable [248], unpredictable, and unfixable—unless assurable world models can be incorporated within LLMs, or within explicit guards. There are many proposals for constructing “guardrails” within or around LLMs. We discuss some of these in the next section on assurance for general-purpose AI but here we will consider external guards and guarding procedures that can be customized to specific applications.
@@
One approach is to develop “workflows” (i.e., sequences of prompts) where the LLM is iteratively asked to critique and improve its previous outputs. A recent paper reports precision around 90% for a workflow that extracts data (as Material, Value, Unit triplets) from materials science papers [187] and similar approaches have been successful in other applications [158]. “Chain of Thought” (COT) [251] is a variant on this that has recently become popular. Another direction exploits the large “context window” (i.e., input) allowed by some recent LLMs to provide a prompt with hundreds of training examples prior to the real query [2] (that paper also proposes ways to automatically generate suitable examples). This is called “in-context learning” [49]; previously, such “fine tuning” required access to the LLM’s training environment and made adjustment to the weights in its neural net. Related to this are applications that provide a substantial input and then ask the LLM to do something with it (e.g., summarize it, or identify its topic). A widely used variant of this is Retrieval Augmented Generation (RAG) [155], where an AI agent retrieves documents from some specified repository and bases its responses on material found there (however, this introduces a new failure mechanism because the retrieval function may fetch incorrect documents).
@@
All these approaches exploit the natural language capability of the LLM, but constrain it to operate on or within the input provided or retrieved, so there is little opportunity to generate or insert “hallucinatory” content extracted from its training corpus. Due to its model-free nature, the LLM may still misinterpret the input and do a wrong thing, but this should also be minimized as a large input can explicitly (via extensive prompts) or implicitly (via a block of text) convey the intended context or model.
3.3 Explanations and Checkable Outputs
@@
Another approach operates by attempting to access the implicit model via the “reasoning” purportedly employed. This is Explainable AI, which aims to deliver reasons for accepting the output from an LLM. Counterfactual explanations, where a system may deliver a response such as “I am declining your loan application but would have approved it if your income was \$5,000 greater, or your deposit was \$2,000 more” [239, 67], may be particularly useful for specific applications as they should be expressed in terms of the policy specified in the prompt and not arbitrary “facts” from the training corpus. A runtime checker could reject decisions whose counterfactual explanations violate the policy. Do note, however, that counterfactual explanations can be manipulated [222], so this approach needs to be employed with caution. However, in some applications it does not matter how a decision and its explanation are produced: they can be considered valid as long as they pass an independent check that incorporates models of the appropriate policy.
@@
A related approach can be applied to AI systems that generate some sort of “design,” such as plans, schedules, software code, or physical artifacts. Even without an explanation, these can often be guarded by analysis tools or simulators for the domain concerned (e.g., see [117] for an elementary application to planning). Again, it does not matter how the proposed design is generated: if it satisfies traditionally engineered and assured verification tools, then it can be considered good.
@@
Automated formalization (or “autoformalization”) is a popular application of this kind that provides an example. The idea is that we have some informal natural language description of the requirements for a computer system, or for a simulation model or for some program code, and the goal is to translate it into a “formal” representation that can be ingested and analyzed, simulated, or executed by some automated verification tool that acts as a guard. A typical approach invites the LLM to generate a formal representation, then uses the verification tool to analyze the result; if this is unsatisfactory, the counterexample, diagnosis, or error report produced by the verification tool is given to the LLM, which is then asked to provide a better formalization. The process iterates until the formal representation passes scrutiny by the verification tool. (This can be seen as a form of “CounterExample-Guided Abstraction Refinement” or CEGAR [61].)
@@
Difficulties may arise with this general approach if the AI system proposes designs that are so original they are outside the capabilities of the analysis or simulation performed by the guarding tool (e.g., a laterally asymmetric airplane may defeat an aerodynamic simulator). A related difficulty arises in domains where there is no reliable means of analysis: for example, generation of strategy for games or real-world scenarios. Applications that generate strategies typically use techniques based on reinforcement learning rather than LLMs and one popular approach, related to mechanisms found in all vertebrate brains [25, Section 6], uses an actor and a critic. The actor generates behaviors and the critic predicts the likelihood that each will lead to a successful outcome. The actor is rewarded by favorable predictions and the critic by their accuracy. Starting with random behaviors and predictions plus some tactic for exploration, the pair will generally converge on optimal behavior, given sufficiently many training examples [123]. This provides assurance for examples in the training set but, as always with ML, tells us little about performance on new examples. One approach is to use the critic as a runtime checker, but the only other automated way to check plausibility prior to commitment seems be use of another (diverse) AI system.
3.4 Diversity